OpenStudy (anonymous):

Fool's problem of the day, Two crispy Arithmetic problems, 1. Five friends- \(A,B,C,D\) and \( E\) - decided to contribute to a charity, \( A \) contributes \( \frac16\) th as much as all the others combined, , \( B \) contributes \( \frac15\) th as much as all the others combined, \( C \) contributes \( \frac14\) th as much as all the others combined, \( D \) contributes \( \frac13\) th as much as all the others combined. If the total contribution of the five is \( \$42,000\) , then how much E pay? Can you find A,B,C,D too (why and why not)? 2. Three solution has acid content \(40 \%, 60 \%\) and \(80 \%\) respectively.When \( x,(x+1)\) and \( (x+2)\) ml respectively are taken from the three and mixed, the resultant is \( 65\% \) acid. What can be said about \(x\)? PS: These are not difficult problems but solving it is the most smartest and least tedious way is somewhat interesting. Good luck!

OpenStudy (anonymous):

x+x+1+x+2=3x+3 40x/100 + (x+1)*60/100 + (x+2)*80/100 = (3x+3)65/100?

OpenStudy (anonymous):

Ishaan94 copied my paper! Gimmeh the medal not him :p

OpenStudy (anonymous):

Me? I do think a question similar to this was posted on openstudy before but I didn't bother to read the thread.

OpenStudy (anonymous):

Haha, im just joking :)

OpenStudy (anonymous):

@Ishaan94: Is it possible to solve it in any other way ;) HINT: Something to do with "Alligation rule".

OpenStudy (anonymous):

E paid $14,000?

OpenStudy (anonymous):

What's Alligation rule? nevermind I'l google it.

OpenStudy (anonymous):

@Ishaan94: E does not pay $14,000

OpenStudy (anonymous):

42000 - 31900 = 10100?

OpenStudy (anonymous):

How you solved it? ;)

OpenStudy (anonymous):

\[\frac{42000 - a}6 = a\]\[\frac{42000 - b}5 = b\]\[\frac{42000 - c}4 = c\]\[\frac{42000 - d}3 = d\] \[e = 42000 - (a+b+c+d)\]I know it's not that creative but it worked.